OpenStudy (anonymous):

How many grams of nitrogen are present in 46.34 g ammonium nitrate?

OpenStudy (anonymous):

mmonium nitrate is NH4NO3 moleculear weight is 80 g / mole nitrogen weight per mole is 28 g / mole % N by weight in a mole = 100 * 28 /80 = 35% 46.34 g amonium nitrate is 46.34*0.35= 16.22 g N

OpenStudy (anonymous):

@lily768 From where you get the 80